OpenStudy (anonymous):

1,16,81,256,625 _____ ?

OpenStudy (goldphenoix):

What's your question?

OpenStudy (anonymous):

whats the next number in the sequence ?

OpenStudy (phi):

try x^4

OpenStudy (mathstudent55):

4th power of numbers

OpenStudy (phi):

notice 1*1*1*1 = 1 2*2*2*2 = 16 3*3*3*3 = 81 and so on....